# Differential Contributions of Machine Learning and Statistical Analysis to Language and Cognitive Sciences

### v1

@@ 3.3.2 Mixed-effect regression analysis
Data collection often involves variables such as the characteristics of speakers, age groups, and gender. Furthermore, the sample sizes available for these studies may be limited, posing challenges when attempting to apply complex models that include a multitude of parameters. Additionally, the assumption of data independence can be problematic. For instance, data may be gathered using quadrats within specific sites, creating a hierarchical data structure with quadrats nested inside the sites. Mixed models are specially designed to manage these types of intricate, structured datasets—even when faced with smaller sample sizes and a large number of variables. Notably, these models also provide the benefit of conserving degrees of freedom, an improvement over traditional linear models in handling such data scenarios. We employed two popular mixed-effect regression models to explore how different factors take effect on durations of words.
@@
The first is Linear Mixed-Effects Regression (LMER) (Bates, , 2010; Kuznetsova et al., , 2017). LMER is a statistical model used to analyze data with both fixed effects, which represent the main variables of interest expected to have a consistent impact across the dataset, and random effects, which account for inherent variations from grouped or nested data sources. This model is particularly useful for dealing with hierarchical structures, such as students within classrooms, or data from repeated measures on the same subjects, allowing for more accurate estimates by accounting for both within-group and between-group variability. LMER helps in understanding complex datasets by modeling the dependencies and structure within the data, making it a powerful tool for robust statistical analysis. The present study employed LMER to explore how independent variables and random variable take effects on the dependent variables. The dependent variable is “word duration”, and other independent variables include word length, word frequency, and the random variables (i.e. Speaker, Age and Sex). The data number (n) in the regression fittings is 262342.
The LMER fittings were summarized as follows, and we made comparison via AIC (Akaike Information Criterion). Control predictors such as word length and word frequency were included, and random variables such as age, sex and speaker could be included. almUnderstanding mixed-effect models is essential for appreciating the significance of these indicators, highlighting the Akaike Information Criterion (AIC) as the preferred tool for model comparison. A smaller AIC indicates better performance, as shown in Table 3.
@@
First, most of factors listed were significant in these fittings. In other words, these factors significantly influenced word duration. Despite this, clearly, the smaller AIC indicates better performance. “lm3” / “lm5””has the smallest AIC, and this suggests that these factors and the random variable have significant effects on word duration. Compared with “lm3” and “lm4”, we find that the factor “Deletion” did significantly contribute to AIC in the model. However, when the random effect “age” or “sex” or both appears, they have no great difference. Through comparison, we found that the random factor “sex” had no significant effect on word duration.
@@
Next, we applied Generalized Additive Mixed Models (GAMM) in analyzing these factors in a similar way. GAMMs are an extension of Generalized Linear Mixed Models (GLMM), incorporating non-linear relationships between the dependent and independent variables through smooth functions (Wood, , 2017). GAMMs allow for both fixed and random effects, accommodating complex variations within hierarchical data structures. The “additive” part of GAMM means that the model expresses the dependent variable as a sum of smooth functions of predictors, along with any random effects and an error term. This flexibility makes GAMMs particularly useful for modeling non-linear trends in data, where the effect of variables is not strictly linear and may vary by group or over time.
@@
We still listed a number of GAMM fittings and made comparison by referring to AIC. The biggest differences between GAMM and LMER is that GAMM could leverage the function s() and interaction smooth te(). The smooth function better gets model fittings for some factors, and the interaction smooth could find the interaction among some given factors. The independent and dependent variables were set similarly as the ones in LMER. The AIC results are shown in Table 4.
@@
Similarly, most of these factors are significant in these cases, which is consistent with the results in LMER. Random effect plays a crucial role in regression analysis (Baayen et al., , 2017). Through the analysis, we found that the random variable “age” is not significant in “t1”. T2 is the best fitting compared with other ones. Comparing “t2” and ”t6”, we could find that the factor “semantic relevance” substantially contributed to the model, namely, the factor “semantic relevance” had a significant impact on word duration.
@@
To further understand the impacts of these factors, we visualized the partial effects of various predictors on word duration, as shown in Fig. 5. We found that word length and “CiteLength” are negatively correlated with word duration, indicating that as these factors increase, the word duration decreases. It is understandable that words with more syllables and letters could be spoken more slowly. Conversely, fast phrase rates and more deletions lead to shorter word durations. Word frequency and semantic relevance exhibit complex effects. When word frequency is low (less than 2), it is positively correlated with word duration; however, when word frequency exceeds 3, its impact turns negative. Similarly, when a word is less semantically related to the context (score less than 2), it negatively affects word duration. In contrast, high semantic relevance (score greater than 2) positively influences word durations.
@@
Figure 5: Partial effects of a given factor on word duration. (Note: The x-axis signifies the metrics, while the y-axis delineates word duration. Each curve visually articulates the relation between a predictor variable and the response variable. A steeper incline on these curves underscores a more robust impact between the predictor and reading speed, whereas gentler slopes imply a less pronounced effect. Moreover, when a curve fluctuates around zero, its effect vanishes. The p-values is smaller than 0.0001 in each plot.)
3.4 Discussion
@@
In summary, we employed machine learning methods such as random forests and SVMs to train on the training dataset, allowing them to identify patterns that were then applied to predict various word durations in the test dataset. In contrast, statistical methods like LMER or GAMM were used primarily to analyze how specific factors affect word duration. These statistical methods do not require training on a testing dataset to determine necessary model parameters, and they are not typically involved in pattern recognition. Consequently, datasets used for statistical analysis do not need to be divided into training and test datasets.
@@
Let us correlate our experiments with the key differences outlined in Table 1 to illustrate their specific applications in research on language and cognitive sciences. First, random forest and SVM were primarily employed to predict varying ranges of word durations in a new dataset. It was observed that both methods achieved a prediction accuracy of 51%. However, correlation analyses and mixed-effects regression demonstrated weaker predictive performance for word durations on new data. Second, machine learning methods often prioritize model accuracy over the interpretability of the extracted features. While the features used in these methods are interpretable, real-world machine learning algorithms focus on feature effectiveness in prediction. In contrast, features (i.e., variables) in statistical analysis should be interpretable from linguistic or cognitive perspectives. In fact, all the factors we selected are highly interpretable. Third, machine learning focuses on how features contribute to the prediction accuracy of word duration. Regression models, however, aim to determine how certain factors affect word duration and the strength of these effects, often within the framework of established statistical theories and constraints. For instance, regression models provide insights into how various factors influence model performance, thereby helping researchers understand the significant impact of these factors on the dependent variable (i.e., word durations). Fourth, statistical analysis addresses the bias and variance of estimates. This focus is crucial as researchers are interested in the distribution of data and their correlations; hence, bias and variance are key considerations.
As mentioned earlier, while machine learning and stastistics are closely related, the choice between employing a machine learning model or a statistical model often hinges on the intended use of the analysis. In some cases, a mixture of machine learning and statistical analysis could be probably used, especially in data-driven research. For instance, in computing the factor semantic relevance, we introduced word embeddings to represent word meanings. Word embedding is a well-known machine learning (specifically, deep learning) technique for deriving word meanings (Mikolov et al., , 2013; Kusner et al., , 2015; Ethayarajh, , 2019), and the algorithm for semantic relevance also incorporates various machine learning methods. However, the computation of semantic relevance remains transparent and interpretable. We employed methods such as Generalized Additive Mixed Models (GAMM) or Linear Mixed Effects Regression (LMER) to investigate how semantic relevance influences word durations. The ultimate goal is to elucidate that semantic relevance is a crucial aspect of human language comprehension; in other words, contextual information plays a vital role when humans produce language. In this sense, GAMM analysis provides interpretable results, further confirming the impact of semantic relevance on the complexity of language production.

### v2

@@ 3.3.1 Correlation
While correlation analysis provides a useful descriptive measure of the relationship between two given factors, it has limitations when exploring complex interactions and relationships. Advanced regression models are necessary to reveal more unique and multifaceted relationships among variables. For instance, as we know, a response variable is supposed to be influenced by a number of factors. When we aim to understand how multiple factors collectively influence this response variable, or when we need to account for random effects, correlation analysis alone is insufficient. Advanced regression models offer other insights: non-linear relationships, random effects, controlling for confounders etc. For example, in language production research, we might use a mixed-effect regression model to examine how factors such as age, gender, word frequency, and speaking rate collectively influence speech duration, while also accounting for random effects of individual speakers and words. This approach would provide a much richer understanding of the complex relationships at play compared to simple correlation analysis.
@@ 3.3.2 Mixed-effect regression analysis
Regression models are fundamental tools in statistical analysis, serving multiple crucial purposes in understanding and quantifying relationships between variables. At their core, these models aim to elucidate how changes in one or more independent variables affect a dependent variable. Regression models enable precise estimation of how much the dependent variable is expected to change when an independent variable is altered, holding other factors constant. Furthermore, regression models are instrumental in controlling for confounding variables, thereby isolating the effects of specific predictors and providing a clearer picture of complex relationships. This function allows researchers and analysts to test hypotheses, and identify the most influential factors in complex systems.
Regression models in statistical analysis encompass a wide range of techniques, from simple linear regression to more complex generalized mixed-effect models. These include ordinary least squares (OLS), logistic regression, polynomial regression, and advanced mixed-effect models like Linear Mixed-Effects Regression (LMER) and Generalized Additive Mixed Models (GAMM). LMER and GAMM are particularly powerful for analyzing hierarchical or nested data structures and handling non-linear relationships, respectively. Simple linear regression and mixed-effect regression models differ significantly in their structure and capabilities. Simple linear regression models the relationship between one dependent variable and one independent variable, assuming a linear relationship and independence of observations. In contrast, mixed-effect regression models can incorporate multiple independent variables and account for both fixed effects (population-level trends) and random effects (group-level variations). This makes them particularly useful for analyzing hierarchical or nested data structures, such as repeated measures or longitudinal studies. Mixed-effect models allow for different intercepts and/or slopes for different groups within the data, providing a more nuanced understanding of relationships that may vary across subgroups. These mixed-effect models offer significant advantages over ordinary regression models by accounting for both fixed and random effects, allowing for the analysis of grouped or longitudinal data, and providing more accurate estimates when dealing with correlated observations.
Table 4 summarizes the steps for implementing and analyzing the relationship between a response variable and independent variables using mixed-effect regression models.
Table 4: Steps for implementing and analyzing statistical regression models
Step Name Description
1 Data preparation Ensure data is clean, properly formatted; check for missing values and outliers; explore data distribution and relationships.
2 Model specification Define response variable; identify the factor of the interest;include the control predictors; determine random effects structure; choose appropriate smoothing functions.
3 Model fitting Use functions like gamm() or bam(); specify model formula; choose distribution family and link function.
4 Model evaluation Check convergence; examine summary statistics; assess model fit; perform model diagnostics.
5 Model comparison Fit alternative models; compare using AIC to understand the contribution of variable (other evaluation criteria: P-values, Confidence Intervals etc.).
6 Interpretation Examine smooth term plots; interpret parametric coefficients; analyze random effects structure.
7 Validation Use cross-validation techniques; check for overfitting.
8 Visualization Create plots of smooth terms and interactions; visualize model predictions and confidence intervals.
9 Reporting Summarize key findings and model performance; present relevant plots and tables.
10 Iteration Refine model based on results and domain knowledge; consider additional predictors or alternative structures if necessary.
1) LMERs:
@@
The first model is Linear Mixed-Effects Regression (LMER) (Bates, , 2010; Kuznetsova et al., , 2017). LMER is a statistical model used to analyze data with both fixed effects, which represent the main variables of interest expected to have a consistent impact across the dataset, and random effects, which account for inherent variations from grouped or nested data sources. This model is particularly useful for dealing with hierarchical structures, such as students within classrooms, or data from repeated measures on the same subjects, allowing for more accurate estimates by accounting for both within-group and between-group variability. LMER helps in understanding complex datasets by modeling the dependencies and structure within the data, making it a powerful tool for robust statistical analysis.
The present study employed LMER to explore how independent variables and random variable take effects on the dependent variable. Our primary question is to explore whether some variables (i.e., semantic relevance) take effect on word duration. This question is clearly distinct from the machine learning models to make classification in Study 1. After knowing the research questions and the factors in the data, we select LMER to build up the setups, which could be followed by Table 4. The dependent variable is “word duration”, and independent variables include word length, word frequency, semantic relevance etc., and the random variables (i.e. Speaker, Age and Sex). The data number (n) in the LMER fittings is 262342.
A typical mixed-effect regression model is consisting of two parts: dependent variable and independent variables. The independent variables are typically composed of three (or four) elements: control predictors, factors of our interest, and random variables, and probably with other predictors. The basic structure of LMER fittings is formulated as follows. Control predictors such as word length and word frequency were included. For instance, longer words tend to have longer durations, whereas more common words typically result in shorter durations (Wright, , 1979). Random variables such as age, sex and speaker could be included. The variable of interest in our study is “semantic relevance,” which is determined by the research question and prior investigations. A significant body of research has demonstrated that other various factors influence word duration in speech. For example, words with more syllables generally have longer durations, a phenomenon closely related to word length. Faster speech rates are associated with shorter word durations, a relationship confirmed in multiple studies (Fujisaki, , 1997; Darling-White and Banks, , 2021; Haley et al., , 2023). Furthermore, words that experience greater phonological simplification often exhibit shorter durations, a trend noted in research on conversational language (Yaruss, , 2000; Baker and Bradlow, , 2009; Waage and Iwarsson, , 2024). These factors, along with control predictors and the predictor of interest, should jointly influence word duration. Therefore, they should be incorporated into statistical regression models. However, it remains unclear whether the semantic relationship between a target word and its context impacts word duration. This knowledge gap is the key motivation for our investigation, and we plan to use mixed-effect regression models to explore this relationship.
[…]
Once the setup of the mixed-effect fitting is complete, we can consider various R packages for implementation, such as “lme4”, “nlme”, and “glmmTMB”. These packages do not require training themselves but provide the necessary tools for fitting and analyzing mixed-effect models. Understanding this is crucial for appreciating the significance of performance indicators, with the Akaike Information Criterion (AIC) emerging as the preferred evaluation criterion for model comparison because AIC is an estimator of prediction error and thereby relative quality of statistical models for a given set of data. The AIC balances model fit against complexity, with smaller values indicating better performance, as illustrated in Table 5. When interpreting AIC, we consider that differences of less than 2 may suggest similar levels of support for the models being compared. Additionally, while AIC is valuable, it should not be used in isolation; other factors such as theoretical grounding, research questions, and result interpretability should also inform model selection.
@@
After running the LMER fittings, we evaluated and interpret the results. First, most of factors listed were significant in these fittings (the threshold of p-value smaller than 0.01). In other words, these factors significantly influenced word duration. The detail of interpreting the fitting result is seen inAppendix C. Despite this, clearly, the smaller AIC indicates better performance. “lm3” / “lm5””has the smallest AIC, and this suggests that these factors and the random variable have significant effects on word duration. Compared with “lm3” and “lm4”, we find that the factor “Deletion” did significantly contribute to AIC in the model. However, when the random effect “age” or “sex” or both appears, they have no great difference. Through comparison, we found that the random factor “sex” had no significant effect on word duration, which is consistent with the feature selection (‘speaker’) in traning ML models. These indicate that features’ selection play a crucial role in both machine learning and stastistical regression analysis.
Additionally, these control predictors are significant across all cases. Previous research has demonstrated the importance of control variables such as “WordLength”, “LogWordFreq”, “CiteLength”, and “PhraseRate”. Our focus is on whether the new variable, semantic relevance, affects word duration. The results from the LMERs consistently show that “semantic relevance” influences word duration, as the variable is significant in all fittings. These findings indicate that the degree of semantic connection between the contextual information and the target word (measured by “semantic relevance”) significantly impacts word duration in spontaneous speech, a form of language production. This suggests that semantic relevance plays a role in language production processes.
2) GAMMs:
@@
Next, we applied Generalized Additive Mixed Models (GAMM) in analyzing these factors in a similar way. GAMMs are an extension of Generalized Linear Mixed Models, incorporating non-linear relationships between the dependent and independent variables through smooth functions (Wood, , 2017). GAMMs allow for both fixed and random effects, accommodating complex variations within hierarchical data structures. The “additive” part of GAMM means that the model expresses the dependent variable as a sum of smooth functions of predictors, along with any random effects and an error term. This flexibility makes GAMMs particularly useful for modeling non-linear trends in data, where the effect of variables is not strictly linear and may vary by group or over time. Using GAMM with the same setups could help cross-verify the results obtained from the LMER fittings. The general mathematical equation for a GAMM is:
$g(E(Y_{i}))=X_{i}\boldsymbol{\beta}+f_{1}(x_{1i})+f_{2}(x_{2i})+\cdots+f_{m}(x_{mi})+Z_{i}\mathbf{b}_{i}+\epsilon_{i}$ (4)
Where $g(\cdot)$ is the link function; $E(Y_{i})$ is the expected value of the response variable for the $i$-th observation; $X_{i}\boldsymbol{\beta}$ represents the parametric fixed effects, where $X_{i}$ is the $i$-th row of the model matrix for parametric terms and $\boldsymbol{\beta}$ is the vector of fixed effect coefficients; $f_{j}(x_{ji})$ are smooth functions of covariates $x_{j}$; $Z_{i}\mathbf{b}_{i}$ represents the random effects, where $Z_{i}$ is the $i$-th row of the random effects model matrix and $\mathbf{b}_{i}$ is the vector of random effects; $\epsilon_{i}$ is the error term.
The smooth functions $f_{j}(\cdot)$ are typically represented using basis expansions, such as splines. The random effects $\mathbf{b}_{i}$ are usually assumed to follow a normal distribution with mean zero and some covariance structure. This equation combines elements of generalized linear models, additive models, and mixed-effects models, allowing for flexible modeling of non-linear relationships while accounting for hierarchical data. Compared with Equation (3), this GAMM equation (4) is basically consistent with the math equation of regression model in machine learning. More details on their differences could be seen in the Discussion Section 4.2.
@@
Implementing a GAMM follows a similar procedure to fitting a LMER fitting, requiring the setup of both dependent and independent variables. The independent variables include control predictors, the factors of our interest and the random variables, which are identical with the ones in LMERs. After running the GAMM fittings, we need to evaluate and interpret the results. We listed a number of GAMM fittings and made comparison by referring to AIC. The biggest differences between GAMM and LMER is that GAMM could employ the function s() and interaction smooth te(). The smooth function better gets model fittings for some factors, and the interaction smooth could find the interaction among some given factors. The independent and dependent variables were set similarly as the ones in LMER. The AIC results are shown in Table 6.
@@
Similarly, most of these factors are significant in these cases, which is consistent with the results in LMER fittings. Random effect plays a crucial role in regression analysis (Baayen et al., , 2017). Interpreting the parameters in GAMM fitting result is seen in Appendix C. Through the analysis, we found that the random variable “age” is not significant in “t1” (p-value is smaller than 0.01), which is consistent with the analysis of LMER fittings and feature selections in training ML models. T2 is the best fitting compared with other ones. Comparing “t2” and ”t6”, we could find that the factor “semantic relevance” substantially contributed to the model, namely, the factor of our interest, “semantic relevance”, had a significant impact on word duration. The result is consistent with the one from the LMER fittings.
@@
To further understand the impacts of these factors, we visualized the partial effects of various predictors on word duration, as shown in Fig. 3. The visualization of partial effects in GAMM results is a valuable tool to make a better understanding of the effects of the included predictors. In addition to p-values and AIC values, the curve trend of partial effect for a given predictor could reveal significant information. The instruction on how to read the curve of partial effects could be seen in the caption of Fig. 3. According to Fig. 3, we found that word length and “CiteLength” are negatively correlated with word duration, indicating that as these factors increase, the word duration decreases. It is understandable that words with more syllables and letters could be spoken more slowly. Conversely, fast phrase rates and more deletions lead to shorter word durations. Word frequency and semantic relevance exhibit complex effects. When word frequency is low (less than 2), it is positively correlated with word duration; however, when word frequency exceeds 3, its impact turns negative. Similarly, when a word is less semantically related to the context (score less than 2), it negatively affects word duration. In contrast, high semantic relevance (score greater than 2) positively influences word durations.
Semantic relevance measure quantifies the degree to which contextual information semantically influences the target word, providing a unique understanding of language processing. The results of our GAMM fittings reveal that semantic relevance significantly impacts word duration in speech, suggesting that words with higher semantic relevance to their context are produced differently than those with lower relevance. This finding opens up exciting avenues for further research into the mechanisms of language production. For instance, researchers could explore how varying levels of semantic relevance affect not only word duration but also other aspects of speech, such as intonation, stress patterns, or articulatory precision. Additionally, investigating how semantic relevance interacts with other linguistic factors, such as word frequency, could provide a more comprehensive model of language production. Recent research has demonstrated the efficacy of semantic relevance in predicting eye movements during reading across multiple languages (Sun, , 2024; Sun, , 2023; Sun et al., , 2024). Given that reading is a quintessential form of language comprehension, this finding, combined with its impact on speech production, establishes semantic relevance as a robust predictor in both language comprehension and production processes. Consequently, this metric appears to play a significant role in human language processing, bridging the gap between how humans comprehend and produce language. The broad applicability of “semantic relevance” suggests its fundamental importance in language processing and opens up new avenues for exploring the intricate cognitive mechanisms.
In summary, the LMERs and GAMMs can cross-verify each other, consistently confirming the effects of these predictors and random variables across the two different models. Our main concern on the influence of “semantic relevance” on word duration could be addressed well: the metric is significant. Further analysis could explore how semantic context affects the processing of target words during language production, which can be examined from both psychological and linguistic perspectives. This underscores the importance of statistical analysis and its interpretations in the fields of language science and cognitive studies.
@@
Figure 3: Partial effects of a given factor on word duration. (Note: The x-axis signifies the metrics, while the y-axis delineates word duration. The plots are arranged in two rows, each containing three graphs. The top row, from left to right, displays the predictors: word length, word frequency, and CiteLength. The bottom row, also from left to right, shows: semantic relevance, phrase rate, and deletions. Each curve for an individual plot visually articulates the relation between a predictor variable and the response variable. A steeper incline on these curves demonstrates a more robust impact between the predictor and word duration, whereas gentler slopes imply a less pronounced effect. Moreover, when a curve fluctuates around zero, its effect vanishes. The p-value is smaller than 0.0001 in each plot.)
4 Discussion
@@
In summary, we employed machine learning methods such as RFs and SVMs to train on the training dataset, allowing them to identify patterns that were then applied to predict the classification of various word durations in the test dataset. In contrast, statistical methods like LMER or GAMM were used primarily to analyze how specific factors affect word duration. These statistical methods do not require training on a testing dataset to determine necessary model parameters, and they are not typically involved in pattern recognition. This section provides a deeper analysis of the differences observed in our experiments, with a focus on the regression models used in both machine learning and statistical analysis. Additionally, we discuss how machine learning and statistical approaches can be applied differently in the context of cognitive science.
4.1 Key differences between machine learning and statistics
Based on Table 3, Table 4, and (a) of Table 1 as well as the Python scripts and R scripts, we overview the differences on implementation of machine learning and regression statistical analysis. The implementation of machine learning models and statistical regression analyses, while sharing some common ground, exhibit distinct approaches, implementations, and emphases. Machine learning typically follows a more iterative, performance-driven process, focusing on predictive accuracy and model optimization.
Machine learning begins with problem definition and data gathering, followed by preprocessing and feature selection. The process then moves through model selection, data splitting, training, and evaluation, with a strong emphasis on tuning and optimizing model performance using metrics like precision, recall, and F1-score. In contrast, statistical regression analysis, particularly when implementing advanced models like GAMMs, adopts a more structured, inference-oriented approach. It starts with meticulous data preparation and exploratory analysis, followed by careful model specification that considers factors of interest, control predictors, and random variables. The analysis then progresses through model fitting, evaluation, and comparison, with a focus on statistical significance like p-value, confidence intervals, and information criteria like AIC. Statistical approaches place greater emphasis on model interpretation, validation of assumptions, and the iterative refinement of models based on both statistical results and domain knowledge. While both methodologies involve data preparation, model fitting, and evaluation, statistical regression analysis typically requires more upfront theoretical consideration and places higher importance on explaining relationships and testing hypotheses, rather than solely optimizing predictive performance.
@@ 4.1 Key differences between machine learning and statistics
Let us correlate our experiments with the key further differences outlined in (b) of Table 1 to illustrate their specific applications in research on language and cognitive sciences. First, random forest and SVM were primarily employed to predict varying ranges of word durations in a new dataset. It was observed that both methods achieved a prediction accuracy of 73% for classifying word durations. However, correlation analyses and mixed-effects regression demonstrated are not required to predict word durations on new data. The reason for this is that ML and statistics are taken to solve different questions: ML is to make predictions for clustering, classification or regression; in contrast, statistics aims to explore the relationship among variables instead of making predictions for new data. Second, machine learning methods often prioritize model accuracy over the interpretability of the extracted features. While the features used in these methods are interpretable, real-world ML algorithms focus on feature effectiveness in prediction. In contrast, features (i.e., variables or factors) in statistical analysis should be interpretable from linguistic or cognitive perspectives. In fact, all the factors we selected are highly interpretable. Third, ML focuses on how features contribute to the prediction accuracy of word duration (e.g. the feature “Speaker”). Statistical regression models, however, aim to determine how certain factors affect word duration and the strength of these effects, often within the framework of established statistical theories and constraints. For instance, GAMMs or LMERs provide insights into how various factors influence model performance, thereby helping researchers understand the significant impact of these factors on the dependent variable (i.e., word durations).
4.2 Differences between regression models used in ML and statistics
Section 2.3 and the LMER/GAMM implementations employed regression models for statistical analysis. It is important to address the distinctions between using regression models in ML and in statistical analysis, as these differences often cause confusion. While regression models in both fields may use the same algorithms, they differ significantly in terms of purpose, methodology, and application. These differences stem from the distinct objectives and methodologies of machine learning and traditional statistical analysis.
ML regression models, such as linear regression, decision trees, or more complex ensemble methods like decison trees in Random Forest, primarily focus on prediction. They excel at capturing complex relationships in large datasets and are often used when the primary goal is to make accurate predictions on new, unseen data. The section2.3 presented an example of using a regression model to predict children’s cognitive development based on early linguistic behaviors. Regression models in ML are well-suited for making future predictions. In this case, regression models are an appropriate choice for such a predictive task. Going back to Equation (3), the training and optimization of the regression model in the context of machine learning aim to find the best set of optimized parameters for this equation. These models typically employ strategies like regularization to prevent overfitting and use cross-validation for model selection and performance evaluation. However, ML regression models often sacrifice interpretability for predictive power. In other words, interpretability is not required for these models when they are applied to predictions.
In contrast, for instance, GAMMs, which fall under the umbrella of statistical analysis, strike a balance between flexibility and interpretability. They extend generalized linear mixed models by incorporating smooth functions to model non-linear relationships between predictors (independent variables) and the response variable (dependent variable). This way, GAMMs are particularly useful when the goal is understanding and interpreting the relationships between variables. The section2.3 mentioned that regression statistical analysis is taken to explore relationships among the children cognitive development and linguistic variables (e.g., syntactic complexity, vocabulary size/diversity, sentence length, speech rate or fluency measures etc.), providing comprehensive insights into the language development trends or characteristics.
Now referring back to Equation (4), we understand that a GAMM needs a process of parameter estimation in this equation, that is, “fitting”, which involves estimating fixed effect coefficients, random effect variances, and smoothing parameters for non-linear functions. Techniques like Maximum Likelihood Estimation (MLE), Restricted Maximum Likelihood (REML), or Penalized Likelihood Estimation are typically employed, often using iterative algorithms such as Penalized Iteratively Reweighted Least Squares (P-IRLS). The goal is to accurately describe the relationships between variables, with a focus on model interpretability. In contrast, a regression model in ML also needs such a process but it is termed “training” (see Equation (3)). The training process in ML often involves minimizing a loss function–commonly via optimization algorithms such as Gradient Descent, Stochastic Gradient Descent (SGD), or more advanced variants like Adam or RMSprop.
Despite their similarities in parameter estimation, GAMMs and machine learning regression models serve fundamentally different purposes. The goal of fitting a GAMM is not to achieve the best predictive performance but to gain insights into the underlying patterns and structures in the data because understanding variable interactions is more important than prediction. Regression models in ML, on the other hand, are primarily employed for making accurate predictions. The training process in ML is focused on optimizing predictive performance, and the goal is to create a model that generalizes well to unseen data. Moreover, the terminology and methodology between GAMMs and machine learning regression models reflect their differing objectives. GAMM fitting focuses on model interpretation, balancing complexity with interpretability through the use of smoothing parameters. Overfitting is controlled by adjusting these parameters, with the aim of understanding the significance of the relationships between variables. Machine learning training, in contrast, emphasizes predictive accuracy. Models are trained to minimize error and maximize generalization, often using advanced optimization and tuning techniques. The complexity of the model is balanced by validation techniques to ensure it performs well on new, unseen data. Additionally, a trained ML model is typically tested on a separate test dataset, whereas a fitted statistical model does not require testing on a split dataset. This difference arises from their distinct objectives: ML models aim to predict outcomes, while statistical models focus on exploring relationships among variables.
GAMMs are often handled by specialized statistical package/software, which may obscure the complexity of the estimation process, while ML models popularize the term “training”, making the process more explicit in modern data science workflows. However, it is important to recognize that both approaches involve sophisticated methods for learning patterns from data, even if their goals and some techniques differ
One key advantage of GAMMs over machine learning regression models is their ability to provide clear visualizations of smooth effects and offer statistical inference on model components. This makes GAMMs particularly valuable in research settings where understanding the nature of relationships and testing specific hypotheses are crucial. GAMMs also handle uncertainty more explicitly, providing standard errors (e.g., AIC) and confidence intervals (e.g., p-value,) for model components.On the other hand, the training process in ML is focused on optimizing predictive performance, and the goal is to create a model that generalizes well to unseen data. Metrics like mean squared error or R-squared on a held-out test set measure this performance. Overall, while their objectives and methods differ, both approaches use iterative procedures to learn patterns from data, reflecting distinct traditions in data analysis.
Additionally, the concept of shrinkage in regression models is the tendency for model performance to decrease when applied to new data. Statistical analysis and ML process shrinkage differently. In statistical approaches, shrinkage can be estimated analytically from the model fit itself, based on the model’s assumptions. This contrasts with machine learning approaches, which typically determine shrinkage empirically using separate test data or cross-validation. The key point is that statistical models can often predict their own generalization performance without requiring additional data, as long as their underlying assumptions hold true. This approach allows for efficient use of data but relies heavily on the validity of the model’s theoretical properties.
In practice, the choice between ML regression models and GAMMs often depends on the specific research questions, the nature of the data, and interpretability. While machine learning models might be preferred for pure prediction tasks or when dealing with extremely complex datasets, GAMMs offer a powerful tool for researchers who need to model non-linear relationships while maintaining interpretability and the ability to perform statistical inference. In some cases, a hybrid approach combining elements of both methodologies might provide the most comprehensive analysis.
4.3 ML vs. statistics in cognitive science